# The dynamics of the Reddit collective action leading to the GameStop short squeeze

### v3

@@ Acknowledgments
Figure S5 displays the Granger Index for various combinations of GME signals at lag 7 (approximately 1 day), both when it is significant (p-value $p<0.1$) and when it is not. The Granger Index in the Reddit-to-Trading Volume direction becomes significant during the Action phase, similar to the analysis at lag 1 reported in the main text. Notably, the magnitude of the Granger Index at lag 7 is higher than at lag 1, and the Reddit-to-Trading Volume direction shows the strongest relationship among all possible combinations. Furthermore, during the Action phase, we observe a significant Granger Index in the Reddit-to-Twitter, as well in the Visibility phase, and Trading Volume-to-Twitter directions.
However, when we consider the Multivariate Vector Autoregression model, we find different results compared to those reported in the main text. Figure S6 displays the different coefficients for each signal in various models at lag 7, both when it is significant (p-value $p<0.1$) and when it is not. We observe that the coefficients are consistently close to zero and not statistically significant, except for the Reddit-to-Twitter coefficient in the upper panel, which becomes significant and positive during the Visibility phase. On the contrary, the coefficients at lag 1 are the same of the analysis reported in the main text. Overall, these results corroborate that the shortest time scales are the more significance in the social-media/market dynamics. Furthermore, with the inclusion of lags, the model’s error diminishes as the information content increases [61].
@@
Supplementary Figure S9: Granger Index capturing the predictive power of a signal on another (using daily log return with a lag of 1 hour) for the following pairs: Trading Volume-to-Reddit (blue), Reddit-to-Trading Volume (red), Trading Volume-to-Twitter (gray) and Reddit-to-Twitter (yellow). The vertical gray dotted line indicates the initiation of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the Tweet by Elon Musk (27 January 2021) that marks the start of the Visibility phase. The triangles correspond to statistical significant values (p-values $<0.1$).
Supplementary Figure S10: Each panel shows the coefficients of the multivariate autoregressive model using daily log return at lag 1 hour predicting Twitter (top), Reddit (middle), and Trading Volume (bottom) activities, from antecedent values of Reddit (orange), Twitter (lightblue), and Trading Volume (blue). The vertical gray dotted line indicates the initiation of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the tweet by Elon Musk (27 January 2021). The triangles refer to statistical significant test (p-value $p<0.1$) within the 15-days window.
@@
Supplementary Materials, Section 5 : Convergent Cross-Mapping
To capture and quantify non-linear coupling relations between variables of the same dynamical systems, we employ a statistical technique known as Convergent Cross Mapping (CCM) [52]. Similar to Granger Test, the key concept behind CCM is to test whether the knowledge of the past states of $x$ variable can improve predictions of $y$ variable’s future states. Unlike Granger Test, CCM does not assume that we can separate the information about $x$ from the rest of the systems, as it is often the case in nonlinear models [64, 65]. Given a time series $x(t)$, we consider the delay vectors $X_{\tau,D}(t)=(x(t),\dots,x(t-(D-1)\tau))$ and for large enough embedding dimension D (larger than the attractor dimensions) the manifold $M_{X}$ identified by $X$ reconstructs the original true manifold $M_{S}$ (or all the topological invariant of the dynamical systems $S$). Given two time series $x(t)$ and $y(t)$, we assume that these are generated from an unknown high-dimensional dynamical systems $S$. Using Takens’ theorem [66, 67, 68], we can construct from each time series $s$ a manifold $M_{s}$ and confront at a given time $t$ the points on one reconstructed manifold $(M_{X})$ with the other $(M_{Y})$. If $X\to Y$, then information from $X$ is stored in $Y$ and gets embedded in $M_{Y}$, thus it follows that we can use $M_{Y}$ to better predict $X$ (the manifold $M_{Y}$ contain the information of both $X$ and $Y$).
$\tilde{Y}_{\tau,D}(t)=(y(t),\dots,y(t-(D-1)\tau),x(t))$ and we perform a neighbor search using the first dimensions to predict last dimension, i.e.$x(t)$. The correlation $\rho_{YX}$ summarises the results of this prediction [52]. CCM requires two criteria to conclude synergistic relation: cross mapping and convergence [52]. Cross mapping is linked to the correlation $\rho$ and the stronger the coupling in one (or both) direction, the higher the value of $\rho$, since the prediction error using one manifold is either lower or equal to that of the other. Convergence refers to improving the cross-mapping with time-series length $T$ (sample size to reconstruct the manifold, referred as library size [52]) and it is related to Takens’ theorem. The more the data, the more precise is the reconstruction of the attractor, resulting in closer nearest neighbors and higher correlation $\rho$.
Supplementary Figure S13: Convergent Cross-Mapping for all the stocks. Correlation $(\rho)$ as a function of length of time series $T$ for optimal lag $\tau$ and optimal dimension $D$ for different combinations of signals: Reddit-to-trading volume (orange, left panel), trading volume-to-Reddit (blue, left panel), trading volume-to-price (purple, center), price-to-trading volume (light green, center), price-to-Reddit (light blue, right panel), Reddit-to-price (pink, right panel). In panel a, BB $(\tau=6,D=5)$, in panel b, AMC $(\tau=6,D=7)$, in panel c, NOK $(\tau=6,D=6)$, and in panel d, GME $(\tau=7,D=5)$.
@@
Supplementary Figure S15: Users shift from NIO to BB, AMC, NOK during January. The plot displays the dynamic shift in user engagement, contrasting the raw number of users “leaving” NIO with the raw number of users “adopting” BB, AMC, NOK over time. “Leaving” users are those who have stopped discussing NIO, while "adopting" users are those initiating discussions for the first time. The vertical gray dotted line indicates the beginning of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the Tweet by Elon Musk (27 January 2021).
Supplementary Figure S16: Raw Number of users talking of stock. Each subpanel represents the number of users discussing GME (red solid line), a stock (same color schemes of the Figure 3 in the main text), and the overlapping about both (dashed lines). Panel a displays the stocks that gained popularity in January: BB (left panel), AMC (central panel) and NOK (right panel). Panel b displays the stocks that are no longer in the top rank: PLTR (left panel), TSLA (central panel) and NIO (right panel). The vertical gray dotted line indicates the beginning of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the Tweet by Elon Musk (27 January 2021).

### v4

@@ 4.3 Granger test and Multivariate Vector Autoregressive Model
Granger Test: p-values and coefficient of the Autoregressive Model
Date Reddit >->- > Trading Volume Trading Volume >->- > Reddit Reddit >->- > Twitter Twitter >->- > Reddit Twitter >->- > Trading Volume Trading Volume >->- > Twitter 2021-01-02 0.909 0.041 0.687 0.742 0.934 0.197 2021-01-03 0.909 0.041 0.687 0.742 0.934 0.197 2021-01-04 0.909 0.041 0.687 0.742 0.934 0.197 2021-01-05 0.606 0.07 0.651 0.804 0.998 0.176 2021-01-06 0.488 0.757 0.846 0.751 0.603 0.225 2021-01-07 0.458 0.815 0.963 0.81 0.62 0.296 2021-01-08 0.606 0.698 0.716 0.734 0.549 0.282 2021-01-09 0.479 0.602 0.742 0.876 0.578 0.266 2021-01-10 0.479 0.602 0.742 0.876 0.578 0.266 2021-01-11 0.479 0.602 0.742 0.876 0.578 0.266 2021-01-12 0.462 0.472 0.833 0.756 0.983 0.3 2021-01-13 0.507 0.855 0.67 0.589 0.987 0.226 2021-01-14 0 0.001 0.703 0.327 0.43 0.636 2021-01-15 0 0.001 0.634 0.531 0.618 0.769 2021-01-16 0 0.003 0.54 0.575 0.811 0.521 2021-01-17 0 0.003 0.54 0.575 0.811 0.521 2021-01-18 0 0.003 0.54 0.575 0.811 0.521 2021-01-19 0 0.006 0.524 0.611 0.884 0.497 2021-01-20 0 0.023 0.571 0.675 0.958 0.603 2021-01-21 0 0.033 0.487 0.714 0.877 0.499 2021-01-22 0 0.03 0.42 0.681 0.93 0.461 2021-01-23 0 0.045 0.515 0.64 0.816 0.649 2021-01-24 0 0.045 0.515 0.64 0.816 0.649 2021-01-25 0 0.045 0.515 0.64 0.816 0.649 2021-01-26 0.002 0.086 0.579 0.722 0.742 0.797 2021-01-27 0.001 0.036 0.769 0.92 0.938 0.91 2021-01-28 0.339 0.663 0.067 0.518 0.741 0.865 2021-01-29 0.26 0.296 0.1 0.487 0.611 0.758 2021-01-30 0.346 0.321 0.084 0.764 0.457 0.946 2021-01-31 0.346 0.321 0.084 0.764 0.457 0.946 2021-02-01 0.346 0.321 0.084 0.764 0.457 0.946 2021-02-02 0.405 0.31 0.126 0.787 0.848 0.571 2021-02-03 0.189 0.474 0.117 0.919 0.369 0.797 2021-02-04 0.342 0.504 0.077 0.967 0.416 0.603 2021-02-05 0.308 0.501 0.127 0.888 0.449 0.714 2021-02-06 0.183 0.602 0.095 0.92 0.482 0.819 2021-02-07 0.183 0.602 0.095 0.92 0.482 0.819 2021-02-08 0.183 0.602 0.095 0.92 0.482 0.819 2021-02-09 0.225 0.739 0.118 0.986 0.287 0.796 2021-02-10 0.29 0.795 0.425 0.842 0.215 0.907 2021-02-11 0.657 0.749 0.999 0.194 0.501 0.938 2021-02-12 0.239 0.783 0.821 0.191 0.805 0.521 2021-02-13 0.408 0.817 0.932 0.302 0.641 0.451
Supplementary Table 11: Results of the Granger test between Reddit, Trading Volume, and Twitter signals (GME). The table reports the p-values (rounded to the third decimal) for the different directions across time windows identified by the Date column. Each row represents a separate test conducted for a specific direction: Reddit anticipating Trading Volume, Trading Volume anticipating Reddit, Reddit anticipating Twitter, Twitter anticipating Reddit, Twitter anticipating Trading Volume, and Trading Volume anticipating Twitter.
Supplementary Figure S8: Coefficients of the autoregressive model capturing the relationship between signals over time. In panel a, the coefficient is computed for the restricted model (with only one signal): Reddit (top) and Trading Volume (bottom). In panel b, the coefficients are computed for the full model predicting Reddit. Each point represents the coefficients over time series spanning from 5 January 2021 to 5 February 2021. The vertical gray dotted line indicates the initiation of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the tweet by Elon Musk (27 January 2021). The triangles refer to statistical significant test (p-value $p<0.1$) within the window.
@@
Figure S9 displays the Granger Index for various combinations of GME signals at lag 7 (approximately 1 day), both when it is significant (p-value $p<0.1$) and when it is not. The Granger Index in the Reddit-to-Trading Volume direction becomes significant during the Action phase, similar to the analysis at lag 1 reported in the main text. Notably, the magnitude of the Granger Index at lag 7 is higher than at lag 1, and the Reddit-to-Trading Volume direction shows the strongest relationship among all possible combinations. Furthermore, we observe a significant Granger Index in the Reddit-to-Twitter direction from the Action and Visibility phases, and in the Trading Volume-to-Twitter direction for the Action phase only.
However, when we consider the Multivariate Vector Autoregression model, we find different results compared to those reported in the main text. Figure S10 displays the different coefficients for each signal in various models at lag 7, both when it is significant (p-value $p<0.1$) and when it is not. We observe that the coefficients are consistently close to zero and not statistically significant, except for the Reddit-to-Twitter coefficient in the upper panel, which becomes significant and positive during the Visibility phase. The coefficients at lag 1 are the same of the analysis reported in the main text. Overall, these results corroborate that the shortest time scales are of more significance in the social-media/market dynamics. Furthermore, with the inclusion of lags, the model’s error diminishes as the information content increases [51].
@@
Supplementary Figure S13: Granger Index capturing the predictive power of a signal on another (using daily log return with a lag of 1 hour) for the following pairs: Trading Volume-to-Reddit (blue), Reddit-to-Trading Volume (red), Trading Volume-to-Twitter (gray) and Reddit-to-Twitter (yellow). The time windows span 15-day, corresponding to 120 points per time series. The vertical gray dotted line indicates the initiation of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the Tweet by Elon Musk (27 January 2021) that marks the start of the Visibility phase. The triangles correspond to statistical significant values (p-values $<0.1$).
Supplementary Figure S14: Each panel shows the coefficients of the multivariate autoregressive model using daily log return at lag 1 hour predicting Twitter (top), Reddit (middle), and Trading Volume (bottom) activities, from antecedent values of Reddit (orange), Twitter (lightblue), and Trading Volume (blue). The time windows span 15-day, corresponding to 120 points per time series. The vertical gray dotted line indicates the initiation of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the tweet by Elon Musk (27 January 2021). The triangles refer to statistical significant test (p-value $p<0.1$) within the 15-days window.
@@
Supplementary Materials, Section 6: Reply speed
Supplementary Figure S16: Daily average (and standard deviation of the mean) of the reply speed, defined as the average time interval between a comment and its direct responses under a Reddit post. The average by post is weighted using the number of comments underneath the post. We apply a 5-day moving average to the time series and consider only weekdays to enhance the chart’s clarity.
@@
Supplementary Figure S17: Raw Number of users talking of stock. Each subpanel represents the number of users discussing GME (red solid line), a stock and both (dashed lines). Panel a displays the stocks that gained popularity in January: BB (left panel), AMC (central panel) and NOK (right panel). Panel b displays the stocks that are no longer in the top rank: PLTR (left panel), TSLA (central panel) and NIO (right panel). The vertical gray dotted line indicates the beginning of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the Tweet by Elon Musk (27 January 2021).
Supplementary Figure S18: Users shift from NIO to BB, AMC, NOK during January. The plot displays the dynamic shift in user engagement, contrasting the raw number of users “leaving” NIO with the raw number of users “adopting” BB, AMC, NOK over time. “Leaving” users are those who have stopped discussing NIO, while ”adopting” users are those initiating discussions for the first time. The vertical gray dotted line indicates the beginning of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the Tweet by Elon Musk (27 January 2021).